OpenStudy (anonymous):

Let f(x) = 24 over 3x-2 Find f(-2).

OpenStudy (anonymous):

You have an F ox x equation, so this means whatever value of x they give you, you are able to plug it into the equation. \[f(x)=24/(3x-2)\]\[f(-2)=24/(3(-2)-2)\]\[f(-2)=?\]

OpenStudy (anonymous):

-3 should solve for his question

OpenStudy (anonymous):